OpenStudy (anonymous):

If f(x) =15x 2 + 10x, find f(-1).

OpenStudy (acxbox22):

plug in -1 for x and solve

OpenStudy (solomonzelman):

is it inverse function, or f of "-1" ?

Nnesha (nnesha):

f(-1) = f(x) = f(-1) so plug -1 in to x

OpenStudy (acxbox22):

f(-1)=15(-1)^2+10(-1) f(-1)=15(1)-10 f(-1)=15-10 f(-1)=????